# Improving Stochastic Cubic Newton with Momentum

### v1

@@ C.4 Convex case
We now assume that the function $f$ is convex and its domain is bounded, more precisely we assume its diameter is bounded by $D$. Under these conditions, we have that :
@@
$\frac{\sqrt{M}}{T}\sum_{t=0}^{T-1}a_{t}\leq\frac{1.2\sqrt{M}}{\sqrt{L}T}\sigma_{g,0}^{3/2}+\frac{16}{L\sqrt{M}T}\sigma_{h,0}^{3}+40\frac{L^{3/8}}{M^{3/8}}\sigma_{g}^{3/2}+\frac{73(46)^{3/2}L^{3/2}}{M^{3}}\tilde{\sigma}_{h}^{3}(\beta)\;,$ (28)
@@
Let $\varepsilon\geq 0$ and let’s assume that for all $t\geq T$ we have $F_{t}\geq\varepsilon$, then:
@@
$\frac{1}{\sqrt{\varepsilon}}$ $\geq\frac{CT}{2}\Big{(}1-\frac{1}{T}\log(\frac{F_{0}}{\varepsilon})\Big{)}-\frac{T}{2\varepsilon^{3/2}\sqrt{M}}\Big{(}\frac{1.2\sqrt{M}}{\sqrt{L}T}\sigma_{g,0}^{3/2}+\frac{16}{L\sqrt{M}T}\sigma_{h,0}^{3}+\mathcal{N}(M)\Big{)},$
where we denoted $\mathcal{N}(M):=40\frac{L^{3/8}}{M^{3/8}}\sigma_{g}^{3/2}+\frac{73(46)^{3/2}L^{3/2}}{M^{3}}\tilde{\sigma}_{h}^{3}(\beta)\;.$ We conclude that:
$T$ $\leq\frac{4}{C}\Big{(}\frac{1}{\sqrt{\varepsilon}}+\frac{C}{2}\log(\frac{F_{0}}{\varepsilon})+\frac{\sigma_{g,0}^{3/2}}{2\varepsilon^{3/2}}+\frac{16\sigma_{h,0}^{3}}{LM\varepsilon^{3/2}}\Big{)},$
for $M$ chosen such that $\frac{\mathcal{N}(M)}{\sqrt{M}C}\leq 8\varepsilon^{3/2}$.
To satisfy the previous condition, we can choose $M=\max\Big{(}100L,1260^{8/3}\frac{D^{4}L\sigma_{g}^{4}}{\varepsilon^{4}},180\frac{\sqrt{DL}\tilde{\sigma}_{h}}{\sqrt{\varepsilon}}\Big{)}$.
Finally, we get:
$T$ $\leq 4032\frac{\sqrt{MD^{3}}}{\sqrt{\varepsilon}}+2\log(\frac{F_{0}}{\varepsilon})+2016\frac{\sqrt{MD^{3}}\sigma_{g,0}^{3/2}}{\varepsilon^{3/2}}+16\times 4032\frac{D^{3/2}\sigma_{h,0}^{3}}{L\sqrt{M}\varepsilon^{3/2}}$
@@
$T=\mathcal{O}\Big{(}\frac{\sqrt{LD^{3}}}{\sqrt{\varepsilon}}+\frac{D^{7/4}L^{1/4}\tilde{\sigma}_{h}^{1/2}}{\varepsilon^{3/4}}+\frac{D^{5/2}L^{1/4}\sigma_{g}}{\varepsilon^{5/2}}+2\log(\frac{F_{0}}{\varepsilon})\\ +(\frac{\sqrt{LD^{3}}}{\varepsilon^{3/2}}+\frac{D^{7/4}L^{1/4}\tilde{\sigma}_{h}}{\varepsilon^{7/4}}+\frac{D^{7/2}L^{1/2}\sigma_{g}^{2}}{\varepsilon^{7/2}})\sigma_{g,0}^{3/2}+\frac{D^{5/4}\sigma_{h,0}^{3}}{L^{5/4}\tilde{\sigma}_{h}^{1/2}\varepsilon^{5/4}}\Big{)}.$
We notice that this upper bound on $T$ is badly affected by the terms proportional to $\sigma_{g,0}^{3/2}$ and $\sigma_{h,0}^{3}$, we remind that $\sigma_{g,0}$ ($\sigma_{h,0}$) represents the variance of the first gradient (Hessian resp.) sample. One simple fix that can improve this rate is to use a large batch to have a good first estimate of the gradient and the Hessian; let this batch have size $b$, then we will have:
$\sigma_{g,0}=\sigma_{g}/\sqrt{b}\;,$
we can, for example, choose $b=\frac{T}{10}$ or any constant other than $10$, then this will only change the rate by a multiplicative factor of $\frac{11}{10}$. By doing this we can show that:
$\frac{\sqrt{MD^{3}}\sigma_{g,0}^{3/2}}{\varepsilon^{3/2}}=\tilde{\mathcal{O}}\Big{(}\frac{L^{1/8}D^{3/8}}{\varepsilon^{9/8}}+\frac{L^{1/16}D^{7/16}\tilde{\sigma}_{h}}{\varepsilon^{19/16}}+\frac{L^{1/8}D^{7/8}\sigma_{g}^{1/2}}{\varepsilon^{13/8}}\Big{)}\sigma_{g}^{3/2}$
and
$\frac{D^{3/2}\sigma_{h,0}^{3}}{L\sqrt{M}\varepsilon^{3/2}}=\tilde{\mathcal{O}}\Big{(}\frac{\sigma_{h}^{3}}{LD^{3/4}M^{2}\varepsilon^{3/4}}\Big{)}.$
Note that all powers of $\varepsilon$ are smaller than $\frac{1}{2}$.
Thus, in this case,
$T=\mathcal{O}\Big{(}\log(\frac{F_{0}}{\varepsilon})+\frac{\sqrt{LD^{3}}}{\sqrt{\varepsilon}}+\frac{D^{7/4}L^{1/4}\tilde{\sigma}_{h}^{1/2}}{\varepsilon^{3/4}}+\frac{D^{5/2}L^{1/4}\sigma_{g}}{\varepsilon^{5/2}}\Big{)}.$

### v2

@@ C.4 Convex case
We assume that the diameter of the domain of $f$ is bounded by $D$. This means that :
@@
$\frac{\sqrt{M}}{T}\sum_{t=0}^{T-1}a_{t}\leq 20\frac{a_{g}^{9/14}\sigma_{g}^{3/2}}{T^{3/7}}+803\frac{a_{h}^{9/5}\sigma_{h}^{3}}{M^{3/2}T^{3/5}}+40\frac{L^{3/8}}{M^{3/8}}\sigma_{g}^{3/2}+\frac{73(46)^{3/2}L^{3/2}}{M^{3}}\tilde{\sigma}_{h}^{3}(\beta)\;,$ (31)
@@
Let $\varepsilon=\underset{0\leq t\leq T-1}{\min}F_{t}$. Thus, for all $t\leq T-1$ we have $F_{t}\geq\varepsilon$, then:
@@
$\frac{1}{\sqrt{\varepsilon}}$ $\geq\frac{CT}{2}\Big{(}1-\frac{1}{T}\log(\frac{F_{0}}{\varepsilon})\Big{)}-\frac{T}{2\varepsilon^{3/2}\sqrt{M}}\mathcal{A}(M),$
where we denoted $\mathcal{A}(M):=20\frac{a_{g}^{9/14}\sigma_{g}^{3/2}}{T^{3/7}}+803\frac{a_{h}^{9/5}\sigma_{h}^{3}}{M^{3/2}T^{3/5}}+40\frac{L^{3/8}}{M^{3/8}}\sigma_{g}^{3/2}+\frac{73(46)^{3/2}L^{3/2}}{M^{3}}\tilde{\sigma}_{h}^{3}(\beta)\;.$ We conclude that:
$\frac{1}{\sqrt{\varepsilon}}$ $+\frac{C}{2}\log(\frac{F_{0}}{\varepsilon})+\frac{T}{2\varepsilon^{3/2}\sqrt{M}}\mathcal{A}(M)\geq\frac{C}{2}T,$
We invert the previous inequality and get:
$\varepsilon$ $\leq\frac{36}{C^{2}T^{2}}+F_{0}e^{-T/3}+\Big{(}\frac{\mathcal{A}(M)}{C\sqrt{M}}\Big{)}^{2/3}$
We conclude that:
$\underset{0\leq t\leq T-1}{\min}F_{t}$ $\leq\mathcal{O}\Big{(}\frac{MD^{3}}{T^{2}}+\frac{LD\tilde{\sigma}_{h}^{2}}{M^{2}}+\frac{L^{1/4}}{M^{1/4}}D\sigma_{g}+F_{0}e^{-T/3}+\frac{a_{g}^{3/7}D\sigma_{g}}{T^{2/7}}+\frac{a_{h}^{6/5}D\sigma_{h}^{2}}{MT^{2/5}}\Big{)}$
We choose $M=\max\Big{(}100L,\frac{L^{1/5}\sigma_{g}^{4/5}T^{8/5}}{D^{8/5}},\frac{L^{1/3}\tilde{\sigma}_{h}^{2/3}}{D^{2/3}},\frac{a_{h}^{3/5}\sigma_{h}T^{4/5}}{D}\Big{)}$.
@@
$\min_{0\leq t\leq T-1}F_{t}=\mathcal{O}\Big{(}\frac{LD^{3}}{T^{2}}+\frac{L^{1/3}D^{7/3}\tilde{\sigma}_{h}^{2/3}}{T^{4/3}}+\frac{L^{1/5}D^{7/5}\sigma_{g}^{4/5}}{T^{2/5}}+F_{0}e^{-T/3}+\frac{a_{h}^{3/5}D^{2}\sigma_{h}}{T^{6/5}}+\frac{a_{g}^{3/7}D\sigma_{g}}{T^{2/7}}\Big{)}.$